# Gueter Josmy Faure1    Jia-Fong Yeh1    Min-Hung Chen2    Hung-Ting Su1 Winston H. Hsu1,3    Shang-Hong Lai4

### v2

@@ Abstract
Existing research often treats long-form videos as extended short videos, leading to several limitations: inadequate capture of long-range dependencies, inefficient processing of redundant information, and failure to extract high-level semantic concepts. To address these issues, we propose a novel approach that more accurately reflects human cognition. [^1] We elaborate on this in section A.6 This paper introduces HERMES: temporal-coHERent long-forM understanding with Episodes and Semantics, a model that simulates episodic memory accumulation to capture action sequences and reinforces them with semantic knowledge dispersed throughout the video. Our work makes two key contributions: First, we develop an Episodic COmpressor (ECO) module that efficiently aggregates crucial representations from micro to semi-macro levels, overcoming the challenge of long-range dependencies. Second, we propose a Semantics reTRiever (SeTR) that enhances these aggregated representations with semantic information by focusing on the broader context, dramatically reducing feature dimensionality while preserving relevant macro-level information. This addresses the issues of redundancy and lack of high-level concept extraction. Extensive experiments demonstrate that HERMES achieves state-of-the-art performance across multiple long-video understanding benchmarks in both zero-shot and fully-supervised settings.
Figure 1: Semantic knowledge and Episodic Memory aggregation illustration: Our Episodic COmpressor (ECO) processes the input video frames and aggregates key episodes, while our Semantics reTRiever (SeTR) identifies high-level semantic cues.
@@ 1 Introduction
Whether we want to programmatically create better video summarization tools, index and retrieve specifics from the vast and ever-expanding array of video content, or improve content moderation and copyright enforcement, we need models that excel at video understanding. Not just short videos with few frames—image models can already handle that—but models that can learn to understand minutes-long videos with thousands of interrelated frames.
@@
Long-form video understanding is challenging for several reasons. First and foremost is the temporal complexity, as it requires handling a large number of frames throughout the video. Second, it requires semantic understanding of high-level concepts as well as the narrative structure. The third challenge is the memory and computational constraints, making it non-trivial to solve the previous two challenges. Attempts to address these issues have been made by researchers who mainly borrow ideas from short videos (Wu & Krahenbuhl, 2021; Miech et al., 2020), which is a more mature area of research encompassing action recognition and video classification, among others, and for which datasets are more abundant. However, we observe that these approaches (using techniques such as max pooling, average pooling, 3D convolutions, and frame concatenation, among others), often do not fully account for the unique characteristics of long videos that distinguish them from a simple concatenation of short video segments. We agree that some short-video modeling ideas, especially those at the spatial level, will also hold for longer ones, but when it comes to long-term modeling, macro-level representations should be extracted efficiently.
@@
To fill this void, we propose temporal-coHERent long-forM understanding with Episodes and Semantics abbreviated as HERMES with two key components. An Episodic COmpressor (ECO) that aggregates key information as we process the video. Imagine walking through a scene, observing and retaining specific episodes such as five people singing a birthday song, then one of them cutting a cake, and so on. This module is designed so that the episodes shape the model’s understanding of the scene sequentially without cluttering it. Complementing the Episodic COmpressor is a SEmantic reTRiever (SeTR), which, as its name suggests, identifies and extracts high-level cues that provide a concise overview of the scene and actions. To revisit the previous example of walking through the scene, one can picture SeTR as having a glimpse at the decorations and instantly comprehending that they are witnessing a birthday party.
@@
  • Through comprehensive evaluation and ablation studies, we validate the efficacy of ECO and SeTR.
  • Our model achieves state-of-the-art performance on four long-form video understanding benchmarks in both zero-shot and fully-supervised settings, notably outperforming the state-of-the-art by a significant 7.3% on LVU (Wu & Krahenbuhl, 2021) and 14.9% on MovieChat-1k (Song et al., 2024).
@@ 2 Problem Setting
Given a long-form video $V={f_{1},f_{2},\ldots,f_{N}}$, where $f_{i}$ represents the $i$-th frame and $N$ is the total number of frames, our objective is to develop a model $\mathbf{M}$ that can efficiently process $V$ and generate an understanding $U$ of its content. This understanding should be capable of answering queries $Q$ or following instructions $I$ related to the video content. Formally, we aim to optimize the function:
@@
Figure 2: HERMES framework overview: We stream through a video window-by-window and extract features using a frozen ViT. Each window feature is processed by the ECO (illustrated in the lower left part of the graph) in an online fashion, discarding redundancies along the way and retaining video episodes that are passed to an episodic Q-Former. The video token bank contains the concatenated features of every window, and SeTR selects only the high-level information to pass to a hierarchical frame-to-sequence Q-Former. The episodic and high-level representations are then concatenated before being fed to the frozen LLM, which outputs a text following the instructions.
3 HERMES
This paper is not about a new LLM or a new way to fine-tune existing LLMs or VLMs. It focuses on leveraging what we know about how humans understand visual scenes to guide the model through the same process. Although this work uses an LLM for autoregressive prediction, the core ideas of episodic memory compression (ECO) and semantic knowledge retrieval (SeTR) can be applied to any model where learning contiguous sequences and high-level representations is required.
@@ 3.2 ECO: Episodic COmpressor
See fig. 2 lower left side for an intuitive illustration of ECO. This module maintains a memory buffer with a maximum number of episodes $E$. Upon receiving a window of $F_{w}$ frame features, we first check whether the existing buffer $\mathcal{M}$ has sufficient bandwidth to support the incoming features. If it does, we simply concatenate them to the buffer; otherwise, we proceed with the compression. At its core, ECO is a distribution process that determines the episode to which a certain frame belongs. It can be summarized as:
=$\mathcal{M}^{\prime}=\begin{cases}\mathcal{M}\oplus F_{w}&\text{if }\|\mathcal{M}\|+\|F_{w}\|\leq E\\ \text{ECO}(\mathcal{M},F_{w})&\text{otherwise}\end{cases}$ (5)
@@
ECO works as follows. As long as concatenating the new window and the existing buffer results in a size greater than $E$, we compute the cosine similarity between each pair of frame features in $\mathcal{M}\oplus F_{w}$. We then iteratively merge the most similar frames until the size constraint $E$ is satisfied.
Specifically,
=$\mathcal{M}^{\prime}=\begin{cases}\mathcal{A}=\mathcal{M}\oplus F_{w}&\\ \text{while }\|\mathcal{A}\|>E:&\\ \quad\left(i^{*},j^{*}\right)=\arg\max_{i\neq j}\frac{\mathcal{A}_{i}\cdot\mathcal{A}_{j}}{\|\mathcal{A}_{i}\|\|\mathcal{A}_{j}\|}&\\ \quad\mathcal{A}_{i^{*}}=\frac{(\mathcal{A}_{i^{*}}+\mathcal{A}_{j^{*}})}{2}&\\ \quad\mathcal{A}=\mathcal{A}\setminus\mathcal{A}_{j^{*}}&\\ \mathcal{M^{\prime}}=\mathcal{A}&\\ \end{cases}$ (6)
[…]
@@ 3.3 Episodic Q-Former
The Episodic Q-Former uses the same architecture as the original Q-Former (Li et al., 2023a) and is loaded with weights pretrained by Dai et al. (2023). However, we introduce a pruning approach similar to ECO to combine and branch queries into episodes over a long video. Given initial queries and instructions, we perform self-attention on these queries and then cross-attention between the queries and the visual representations from ECO ($\mathcal{M}^{\prime}$). The enhanced queries then undergo an ECO-like process, where we iteratively merge similar queries across windows, effectively forming episodes of high information density. Such a process helps in keeping a constant number of queries throughout the video. The following equation summarizes the process,
$Q=\text{$\mathfrak{q}$ECO}\left(\text{CA}\left(\text{SA}(Q_{0}),\mathcal{M}^{\prime}\right)\right)$ (7)
@@ 3.4 SeTR: Semantics reTRiever
To capture higher-level semantic information from our video features, we introduce SeTR (Semantics reTRiever). SeTR is designed to identify and consolidate important high-level information that may be scattered (contiguously or not) throughout the video. Given a video feature tensor $F\in\mathbb{R}^{B\times N\times T\times C}$, where $B$ is the batch size, $N$ is the number of frames, $T$ is the number of tokens per frame and $C$ is the channel dimension, SeTR operates as follows: we first normalize $F$ to ensure consistent scaling across features. Second, we apply a stride of $k$ to create two groups, group $X$ containing every $k$-th frame, resulting in $\frac{N}{k}$ frames and group $Y$ with the remaining $N-\frac{N}{k}$ frames. Third, we calculate dot product similarity scores between frames in $X$ and $Y$. Finally, for each frame in $Y$, we merge it with its most similar frame in $X$, based on the computed scores by taking their mean.
This process effectively reduces the number of frames from $N$ to $\frac{N}{k}$, consolidating semantic information while maintaining the most relevant features across the video timeline. The resulting semantic representations are denoted as $F^{\prime}\in\mathbb{R}^{B\times\frac{N}{k}\times T\times C}$. We evaluate the effectiveness of this approach in section 4.3. While Bolya et al. (2022) have explored token reduction in vision transformers, their approach and objectives differ significantly from ours. Their method focuses on minor token reductions within individual frames, specifically between different layers of a Vision Transformer. In contrast, SeTR takes a fundamentally different approach by aggressively eliminating redundancies at the frame level.
@@ 3.5 Hierarchical QFormer
Following our Semantics reTRiever, is a hierarchical Q-Former composed of a frame Q-Former ($fQFormer$), a frame-to-sequence adapter and a video Q-Former ($vQFormer$). The frame Q-Former enhances each semantic piece of information, independently of the others, and the video Q-Former consolidates them. The resulting query $Q_{sem}\in\mathbb{R}^{B\times q\times C^{\prime}}$ contains semantic representations of the video.
@@ 3.6 From Representations to Natural Language
After obtaining the episodic representations $Q$ and the semantic representations $Q_{sem}$, we prepare them for input into a Large Language Model (LLM). Specifically, we concatenate $Q$ and $Q_{sem}$ to form a unified representation vector. This concatenated vector is then projected into the input embedding space of the LLM using a learned linear transformation. In our implementation, we utilize a Vicuna-7B model (Chiang et al., 2023) as LLM. The model, conditioned on this projected representation and guided by task-specific instructions, generates the requested natural language output. This approach allows us to leverage the LLM’s pretrained knowledge and language generation capabilities while incorporating our task-specific episodic and semantic information. The process can be summarized by the following equation:
@@ 4 Experiments
4.1 Datasets
Our framework is designed to be versatile, accommodating tasks that involve either video inputs or a combination of video and written instructions, producing text-based outputs. We evaluated our approach on two primary tasks: long-form video classification and long-form video question answering.
For long-form video classification, we utilize three diverse datasets. The first, LVU (Wu & Krahenbuhl, 2021), focuses on movie content, offering a rich source of narrative and thematic video data. The second, Breakfast (Tang et al., 2019), consists of instructional videos that emphasize procedural and step-by-step understanding. Lastly, COIN (Kuehne et al., 2014) is another instructional video dataset, but it covers a broader range of procedural activities compared to Breakfast.
For long-form video question answering, we employ the MovieChat-1k dataset (Song et al., 2024). This dataset allows us to test our framework under both zero-shot and fully supervised scenarios.
@@ 4.2 Quantitative Results
For the LVU, Breakfast, and COIN datasets, we report top-1 classification accuracy. For the MovieChat-1k dataset, we follow the evaluation protocol developed by Maaz et al. (2023), employing GPT-3.5-turbo (Brown et al., 2020) to assess both accuracy (in %) and answer quality scores (from 1 to 5).
In addition to the zero-shot results on MovieChat-1k, we also include metrics for our model finetuned on the MovieChat-1k training set for one epoch. These trained results are provided for reference but are de-emphasized in the table. Note that for the zero-shot evaluation, we use a flat video Q-Former instead of a hierarchical Q-Former since the later would introduce untrained weights.
Table 1: SOTA Comparison on the LVU Dataset: The table presents Top-1 accuracy for various models. The highest score is highlighted in bold, and the second highest is underlined. Unlike the minor incremental improvements observed among other methods, our model demonstrates a significant performance leap, outperforming its nearest competitor by an average of 7.3%.

### v3

# Gueter Josmy Faure1    Min-Hung Chen2    Jia-Fong Yeh1    Hung-Ting Su1 Shang-Hong Lai4    Winston H. Hsu1,3

@@ Abstract
Existing research often treats long-form videos as extended short videos, leading to several limitations: inadequate capture of long-range dependencies, inefficient processing of redundant information, and failure to extract high-level semantic concepts. To address these issues, we propose a novel approach that more accurately reflects human cognition. [^1] We elaborate on this in Section A.7 This paper introduces HERMES: temporal-coHERent long-forM understanding with Episodes and Semantics, a model that simulates episodic memory accumulation to capture action sequences and reinforces them with semantic knowledge dispersed throughout the video. Our work makes two key contributions: First, we develop an Episodic COmpressor (ECO) module that efficiently aggregates crucial representations from micro to semi-macro levels, overcoming the challenge of long-range dependencies. Second, we propose a Semantics reTRiever (SeTR) that enhances these aggregated representations with semantic information by focusing on the broader context, dramatically reducing feature dimensionality while preserving relevant macro-level information. This addresses the issues of redundancy and lack of high-level concept extraction. Extensive experiments demonstrate that HERMES achieves state-of-the-art performance across multiple long-video understanding benchmarks in both zero-shot and fully-supervised settings. Our code will be made public.
Figure 1: Semantic Knowledge and Episodic Memory Aggregation: Our Episodic COmpressor (ECO) processes input video frames and aggregates key episodes: (I) A cosmic setting with a planet and starfield, (II) A narrator explaining the scene, and (III) An aircraft viewed from inside and out. Concurrently, our Semantics reTRiever (SeTR) identifies high-level semantic cues throughout the video, including: (I) The theme of space exploration and (II) Human interaction with technology. This dual-level approach enables comprehensive video understanding by capturing both specific events and overarching concepts.
@@ 1 Introduction
Video understanding reflects how humans perceive the world through one of our most essential senses, sight, and drives a wide range of visual and multimodal applications. Whether we want to create better video summarization tools, index and retrieve specifics from the vast and ever-expanding array of video content, or improve content moderation and copyright enforcement, we need models that excel at video understanding. This requirement extends beyond short videos with few frames—a task that image models can already handle adequately—to encompass the analysis of extended video content spanning minutes and comprising thousands of interrelated frames.
@@
Long-form video understanding is challenging for several reasons. First and foremost is the temporal complexity, as it requires handling a large number of frames throughout the video. Second, it requires a semantic understanding of high-level concepts as well as the narrative structure. The third challenge is the memory and computational constraints, making it non-trivial to solve the previous two challenges. Attempts to address these issues have been made by researchers who mainly borrow ideas from short videos (Wu & Krahenbuhl, 2021; Miech et al., 2020), which is a more mature area of research encompassing action recognition and video classification, among others, and for which datasets are more abundant. However, these approaches, which adopt techniques such as pooling (Faure et al., 2023), or 3D convolutions (Feichtenhofer et al., 2019), often do not fully account for the unique characteristics of long videos that distinguish them from a simple concatenation of short video segments. Some ideas about short-video modeling, especially for those at the spatial level, may also hold for longer ones, but when it comes to long-term modeling, macro-level representations should be extracted efficiently.
@@
In video understanding, we can distinguish between two types of information: episodic and semantic. Episodic information refers to specific, sequential events that occur in a video, while semantic information encompasses overarching themes and concepts. To illustrate, imagine walking through a scene at a birthday party. Episodic information might include observing five people singing a birthday song, followed by one person cutting a cake. These are specific, time-bound events. In contrast, semantic information might involve recognizing decorations scattered throughout the scene, instantly comprehending that you’re witnessing a birthday party. This high-level understanding provides a concise overview of the scene and actions, transcending specific moments. Building on these concepts, we propose temporal-coHERent long-forM understanding with Episodes and Semantics (HERMES), a model designed to capture both episodic and semantic information in long-form videos. HERMES  comprises two key components: ECO  and SeTR. The Episodic COmpressor (ECO) aggregates key contiguous information as we process the video, shaping the model’s understanding of the scene sequentially without cluttering it. Complementing this, the SEmantic reTRiever (SeTR) identifies and extracts high-level cues that provide a concise overview of the scene and actions. HERMES  achieves state-of-the-art performance on four long-form video understanding benchmarks in both zero-shot and fully-supervised settings, notably outperforming the state-of-the-art by a significant 7.3% on LVU(Wu & Krahenbuhl, 2021) and 14.9% on MovieChat-1k (Song et al., 2024).
@@
  • We address the challenge of efficiently processing and understanding long-form videos, a problem that has been hindered by computational constraints and the difficulty of capturing complex temporal dynamics over extended durations.
@@
Through comprehensive evaluation across multiple benchmarks and detailed ablation studies, we validate the effectiveness of ECO and SeTR, and demonstrate their complementary roles in enhancing long-form video understanding
@@ 2 Problem Setting
Given a long video $V=\{f_{1},f_{2},\ldots,f_{N}\}$, where $f_{i}$ represents the $i$-th frame and $N$ is the total number of frames, our objective is to develop a model $\mathbf{M}$ that can efficiently process $V$ and construct an internal understanding $U$ of its content. This understanding should enable the model to answer queries $Q$ or follow instructions $I$ related to the video content. Formally, we aim to optimize the function:
@@
Figure 2: HERMES framework overview: We stream through a video window-by-window and extract features using a frozen ViT. Each window feature is processed by the Episodic COmpressor (ECO) in an online fashion, discarding redundancies along the way and retaining video episodes that are passed to an episodic Q-Former. The video token bank contains the concatenated features of every window, and SeTR selects only the high-level information to pass to a hierarchical frame-to-sequence Q-Former. The episodic and high-level representations are then concatenated before being fed to the frozen LLM, which outputs a text following the instructions.
3 Proposed Framework: HERMES
This paper is not about a new LLM or a new way to fine-tune existing LLMs or VLMs. It focuses on leveraging what we know about how humans understand visual scenes to guide the model through the same process. Although this work uses an LLM for autoregressive prediction, the core ideas of episodic memory compression (ECO) and semantic knowledge retrieval (SeTR) can be applied to other models where learning contiguous sequences and high-level representations is advantageous.
@@ 3.1 Window Encoder
Algorithm 1 ECO: Episodic COmpressor
1:$\mathcal{A}\leftarrow\mathcal{M}\oplus F_{w}$
2:while $\|\mathcal{A}\|>E$ do
3:     $(i^{*},j^{*})\leftarrow\arg\max_{i\neq j}\frac{\mathcal{A}_{i}\cdot\mathcal{A}_{j}}{\|\mathcal{A}_{i}\|\|\mathcal{A}_{j}\|}$
4:     $\mathcal{A}_{i^{*}}\leftarrow\frac{(\mathcal{A}_{i^{*}}+\mathcal{A}_{j^{*}})}{2}$
5:     $\mathcal{A}\leftarrow\mathcal{A}\setminus\mathcal{A}_{j^{*}}$
6:end while
7:$\mathcal{M}\leftarrow\mathcal{A}$
@@ 3.2 ECO: Episodic COmpressor
The proposed Episodic COmpressor (ECO)  aggregates the video frames into episodes. This module maintains a memory buffer with a maximum number of episodes $E$. Upon receiving a window of $F_{w}$ frame features, we first check whether the buffer $\mathcal{M}$ has sufficient bandwidth to support the incoming features. If it does, we simply concatenate them to the buffer; otherwise, we proceed with the compression. At its core, ECO is a distribution process that determines the episode to which a certain frame belongs. It can be summarized as:
=$\mathcal{M}=\begin{cases}\mathcal{M}\oplus F_{w}&\text{if }\|\mathcal{M}\|+\|F_{w}\|\leq E\\ \text{ECO}(\mathcal{M},F_{w})&\text{otherwise}\end{cases}$ (5)
@@
ECO works as Algorithm 1. As long as concatenating the new window and the buffer results in a size greater than $E$, we compute the cosine similarity between each pair of frame features in $\mathcal{M}\oplus F_{w}$. We then iteratively merge the most similar frames until the size constraint $E$ is satisfied. Specifically,
where $\mathcal{M}$ is the existing buffer, $F_{w}$ represents the incoming window of $w$ frame features, $\mathcal{A}$ is the concatenated buffer and new window, and $\|\mathcal{A}\|$ is the size of $\mathcal{A}$. To summarize Algorithm 1, $\frac{\mathcal{A}_{i}\cdot\mathcal{A}_{j}}{\|\mathcal{A}_{i}\|\|\mathcal{A}_{j}\|}$ computes the cosine similarity between frame features $\mathcal{A}_{i}$ and $\mathcal{A}_{j}$, $\arg\max_{i\neq j}$ finds the pair of frames with the highest cosine similarity, $\frac{(\mathcal{A}_{i^{*}}+\mathcal{A}_{j^{*}})}{2}$ combines the most similar frames, and $\mathcal{A}\setminus\mathcal{A}_{j^{*}}$ removes the frame $\mathcal{A}_{j^{*}}$ from $\mathcal{A}$ after merging. The process repeats until the size of $\mathcal{A}$ is within the maximum allowed episodes $E$.
@@ 3.3 Episodic Q-Former
The Episodic Q-Former uses the same architecture as the original Q-Former (Li et al., 2023a) and is loaded with weights pretrained by Dai et al. (2023). However, we insert ECO as a pruning module within the Q-Former to combine and branch queries into episodes over the long video. Given initial queries and instructions, we perform self-attention on these queries and then cross-attention between the queries and the visual representations $\mathcal{M}$. The enhanced queries then undergo an ECO-like process, where we iteratively merge similar queries across video windows, effectively forming video query episodes of high information density. The following equation summarizes the process,
$Q=\text{ECO}_{\mathfrak{q}}\left(\text{CA}\left(\text{SA}(Q_{0}),\mathcal{M}\right)\right)$ (6)
@@ 3.4 SeTR: Semantics reTRiever
To complement ECO and capture higher-level semantic information from the video, we develop a Semantics reTRiever (SeTR). SeTR is designed to identify and consolidate important high-level information that may be scattered (contiguously or not) throughout the video. Given a video feature tensor $F\in\mathbb{R}^{B\times N\times T\times C}$, where $B$ is the batch size, $N$ the number of frames, $T$ the number of tokens per frame and $C$ the channel dimension, SeTR operates as follows: we first normalize $F$ to ensure consistent scaling across features. Second, we apply a stride of $k$ to create two groups, group $X$ containing every $k$-th frame, resulting in $\frac{N}{k}$ frames and group $Y$ with the remaining $N-\frac{N}{k}$ frames. Third, we calculate dot product similarity scores between frames in $X$ and $Y$. Finally, for each frame in $Y$, we merge it with its most similar frame in $X$, based on the computed scores by taking their mean.
This process effectively reduces the number of frames from $N$ to $\frac{N}{k}$, consolidating semantic information while maintaining the most relevant features across the video timeline. The resulting semantic representations are denoted as $F^{\prime}\in\mathbb{R}^{B\times\frac{N}{k}\times T\times C}$. We evaluate the effectiveness of this approach in Section 4.3. While ToMe (Bolya et al., 2022) have explored token reduction in vision transformers, their approach and objectives differ significantly from ours. Their method focuses on minor token reductions within individual frames, specifically between different layers of a Vision Transformer. In contrast, SeTR retains the most salient frames while significantly reducing redundancies.
@@ 3.5 Hierarchical QFormer
Following our SeTR, is a hierarchical Q-Former composed of a frame Q-Former ($fQFormer$), a frame-to-sequence adapter and a video Q-Former ($vQFormer$). The frame Q-Former enhances each semantic piece of information, independently of the others, and the video Q-Former consolidates them. The resulting query $Q_{sem}\in\mathbb{R}^{B\times q\times C^{\prime}}$ contains the semantic representations of the entire video.
@@ 3.6 From Representations to Natural Language
After obtaining the episodic representations $Q$ and the semantic representations $Q_{sem}$, we prepare them for input into a Large Language Model (LLM). Specifically, we concatenate $Q$ and $Q_{sem}$ to form a unified representation vector. This concatenated vector is then projected into the input embedding space of the LLM using a learned linear transformation. In our implementation, we utilize a Vicuna-7B model (Chiang et al., 2023) as LLM. The model, conditioned on this projected representation and guided by task-specific instructions, generates the requested natural language output. This approach allows us to leverage the LLM’s pretrained knowledge and language generation capabilities while incorporating our task-specific episodic and semantic information. The process is summarized by the following equation:
@@ 4 Experiments
4.1 Datasets and Evaluation Metrics
We evaluate our approach on two primary tasks: long-form video classification and long-form video question answering.
For long-form video classification, we utilize three diverse datasets. The first, LVU (Wu & Krahenbuhl, 2021), focuses on movie content, offering a rich source of narrative and thematic video data. The second, Breakfast (Tang et al., 2019), consists of instructional videos that emphasize procedural and step-by-step understanding. Lastly, COIN (Kuehne et al., 2014) is another instructional video dataset, but it covers a broader range of procedural activities compared to Breakfast. We report top-1 classification accuracy on these datasets.
For long-form video question answering, we employ the MovieChat-1k dataset (Song et al., 2024) and report both zero-shot and fully-supervised results. As evaluation metrics, we follow the evaluation protocol developed by Maaz et al. (2023), employing GPT-3.5-turbo (Brown et al., 2020) to assess both accuracy and answer quality score.
@@ 4.2 Quantitative Results
Table 1: SOTA Comparison on the LVU Dataset: The table presents Top-1 accuracy for various models. Unlike the minor incremental improvements observed among other methods, our model demonstrates a significant performance leap, outperforming its nearest competitor by 7.3%. The highest score is highlighted in bold, and the second highest is underlined.